OpenStudy (anonymous):

Howdo you test for convergence and divergence and what does it mean to be conditional or absolute?

OpenStudy (anonymous):

conditional or absolute convergence exist only in alternate series

OpenStudy (anonymous):

There are many convergence tests.

OpenStudy (anonymous):

A series converges absolutely when \[\sum_{}^{}|a _{k}|\] converges.

OpenStudy (anonymous):

1) divergent test 2) integral test 3) root test 4) ratio test

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty}n/(2n+3)\] for example

OpenStudy (anonymous):

It diverges, read about the harmonic series.

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty}\frac{n}{2n+3}=\frac{1}{2}\] so the series diverges.

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty}a _{n}=0\] is a necessary property.

OpenStudy (anonymous):

Depending on the series you will need to use different convergence tests. You should get a text that discusses it in detail.